A garden is in the shape of a square region. Area of the garden is 40000 square meters. How much distance a man travels if he takes one round of the garden
Answer by stanbon(75887)

A garden is in the shape of a square region. Area of the garden is 40000 square meters. How much distance a man travels if he takes one round of the garden
----
Area formula: A = (side)^2
---
Solve for side::
40000 = s^2
s = sqrt(40000) = 200
---
Perimeter = 4s
Ans: Perimeter = 4*200 = 800 meters
--------------
Cheers,
Stan H.
==================